ganeshie8 (ganeshie8):

show that \(\large \frac{10^{12n-6}-1}{99} \) is not divisible by \(101\) for all \(n \in \mathbb{Z^{+}}\)

ganeshie8 (ganeshie8):

In other words show that below equation is not solvable in integers \[\large 10^{12n-6}-1 = 99\cdot 101k\]

OpenStudy (mathmath333):

is their a property like \(\dfrac{a}{b} (mod~c)=\dfrac{a (mod~c)}{b (mod~c)} \)

ganeshie8 (ganeshie8):

thats a good start! lets consider an example \[\dfrac{6}{3} (mod~3) \stackrel{?}{=}\dfrac{6 (mod~3)}{2 (mod~3)} \]

OpenStudy (mathmath333):

thats invalid

ganeshie8 (ganeshie8):

\[2(mod~3) \stackrel{?}{=}\dfrac{0}{2 (mod~3)} \]

ganeshie8 (ganeshie8):

so such a rule is not valid

ganeshie8 (ganeshie8):

however this is true : \[a\pmod n \times b\pmod n = ab \pmod n\]

ganeshie8 (ganeshie8):

another example to see the mods are not friendly with division : \[\dfrac{6}{4} (mod~3) \stackrel{?}{=}\dfrac{6 (mod~3)}{4 (mod~3)}\]

hartnn (hartnn):

\(10^{12n-6}-1\) is indded divisible by 99, any way to show \((10^{12n-6}-1)/99\) it in terms of 'n' without any denominator... bdw, whats the general way to show an integer is not divisible by another integer...? contradiction , or something else in mod arithmetic or what u posted in 1st reply?

hartnn (hartnn):

tp prove \(10^{12n-6}-1 ~ mod ~9999 \ne 0 \)

ganeshie8 (ganeshie8):

yes \(\dfrac{10^{12n-6}-1}{99}\) gives a nice pattern of numbers : 10101 101010101010101 etc..

OpenStudy (anonymous):

10^2=1 mod 99 10^12n-6=1 mod 99 10^(12n-6) -1=0 md 99 this how u show its integer

hartnn (hartnn):

\([10^{6n-3}-1~ mod ~9999] \times [10^{6n-3}+1~ mod ~9999] \ne 0 \)

ganeshie8 (ganeshie8):

we want to show \(10^{12n-6}-1 \not \equiv 0 \pmod {99 \cdot 101}\)

ganeshie8 (ganeshie8):

that looks interesting, you factored xD

hartnn (hartnn):

just to use \(a\pmod n \times b\pmod n = ab \pmod n \) :P

OpenStudy (anonymous):

i have some idea for it :P will come back after lunch ...

hartnn (hartnn):

\(10^{(6n-3)}-1 |9 \\ \implies 10^{(6n-3)}-1 ~mod~ 1111\) should be shown non-0

ganeshie8 (ganeshie8):

that works i think but we also need to show that \(10^{6n-3}+1 \pmod{1111}\) is not 0

hartnn (hartnn):

i wasn't sure that ...+1 lso divides 9 or not

ganeshie8 (ganeshie8):

Oh I see yeah we still need to show \(10^{6n-3}+1 \pmod{9999}\) is not 0

hartnn (hartnn):

YES, does that ...+1 divide 11 ?

OpenStudy (anonymous):

\(\begin{align*} 10^{12n-6} -1 &=(10^{6n-3} -1) (10^{6n-1} +1) \\ &=(10^{3(2n-1)} -1) (10^{3(2n-1)} +1) \\ &= ((-10)-1 )(-10 +1)\mod 101\\ &= -11*-9= -2 \mod 101 \end{align*} \)

hartnn (hartnn):

what about the 99 lol

OpenStudy (anonymous):

hehe was just spreading ideas :D

OpenStudy (anonymous):

since it does not devide 101 then it does not devide k101 for any k let k be 99 so :P

hartnn (hartnn):

how is \(\Large 10^{3(2n-1)}~mod ~101 = -10\) ?

ganeshie8 (ganeshie8):

that looks perfect ! @Marki

ganeshie8 (ganeshie8):

since it does not devide 101 then it does not devide k101 for any k let k be 99 \(\blacksquare\)

ganeshie8 (ganeshie8):

thats it! thats the end of proof xD

OpenStudy (mathmath333):

is the answer \(\Large \equiv 1 (mod ~~101) \)

OpenStudy (mathmath333):

over integers+

OpenStudy (anonymous):

10^3=(-10) mod 101 =10* (-1) mod 101 2n-1 is odd so (-1)^3(2n-1 ) = -1 mod 101 now we need to show 10^3=10 ^3k mod b let k be 2n-1 thus we have 10^3(2n-1)=(-10) mod 101

OpenStudy (anonymous):

idk if i made some typos but @mathmath333 i showed it for 101 only not for both

OpenStudy (anonymous):

for both i need Diophantine to find mod and since @ganeshie8 asked only for showing its not divisible without finding exact mod then there is no need

OpenStudy (anonymous):

i havn't read ur comments , i just post my idea : @ganeshie8 , @Marki it's just like the idea for 17 : Let suppose that it's divisible by 101,then it can be written \( a=101t \). \( a \equiv 1 ~ (mod ~ 10) \) so \( t \equiv 1 ~ (mod ~ 10) \) too (it can be proven).then \(t=10k+1 \) therefore \(a=1010k + 101 \).

OpenStudy (anonymous):

what if k=5 then a=101000101 ( which is not our sequence )

ganeshie8 (ganeshie8):

@PFEH.1999 we can conclude below from Marki's first reply : \[10^{12n-6}-1 \equiv -2 \pmod {101}\]

ganeshie8 (ganeshie8):

since \(10^{12n-6}-1\) is not divisible by \(101\) itself, it cannot be divisible by \( 99\cdot 101\) also end of story!

hartnn (hartnn):

so it doesn't matter what the denominator is :P unless it makes the fraction as integer

OpenStudy (anonymous):

@ganeshie8 i said i havn't been following ur ideas XD

ganeshie8 (ganeshie8):

Exactly! if a number is not divisible by "2", then we can say that it is also not divisible by any multiple of 2

OpenStudy (anonymous):

so it doesn't matter what the denominator is :P unless it makes the fraction as integer hmm @hartnn at least i showed it gives integer :P

OpenStudy (mathmath333):

continuing marki's solution \(\large \tt \begin{align} \color{black}{\dfrac{1}{99}\dfrac{-2}{101}\\~\\ =\dfrac{1}{99}\times \dfrac{-2}{101}\\~\\ =\dfrac{1}{99} \times \dfrac{-101+99}{101}\\~\\ =\dfrac{1}{99}(99)\\~\\ \Large \equiv 1 }\end{align}\)

hartnn (hartnn):

i am yet to digest why \(\Large 10^{3(2n-1)}~mod ~101 = -10\)

ganeshie8 (ganeshie8):

Ahh that works @PFEH.1999 i was stuck with analyzin g Marki's solution

OpenStudy (anonymous):

ok i'll show it again with complicating it

OpenStudy (mathmath333):

@ganeshie8 is my soln valid

ganeshie8 (ganeshie8):

one sec, im going thru..

ganeshie8 (ganeshie8):

first let me answer @hartnn 's question on why \(10^{3(2n-1)}~\mod ~101 = -10\)

ganeshie8 (ganeshie8):

\[10^{3(2n-1)} \equiv 1000^{2n-1} \equiv (101\times 10 - 10)^{2n-1}\equiv (-10)^{2n-1} \pmod{101}\]

OpenStudy (anonymous):

\(\large \begin{align*} 10^{12n-6}-1 &= (10^2)^{6n-3}-1\\ &= (-1)^{3(2n-1)}-1 (\mod 101) \text{ but 2n-1 is odd and 3 }\\ &= (-1)^k-1 (\mod 101) \text{ for some odd k}\\ &= -1 -1 (\mod 101)\\ &= -2 (\mod 101) \end{align*}\)

OpenStudy (anonymous):

this makes it more simple

ganeshie8 (ganeshie8):

thats kinda cute :D

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

@Marki , nice ;) u are very good at mod properties

OpenStudy (anonymous):

yeah ik im pure math genius :D

OpenStudy (mathmath333):

\(\large \tt \begin{align} \color{black}{\implies \frac{10^{12n-6}-1}{101}\\~\\ \implies \frac{100^{6n-3}-1}{101}\\~\\ \implies \frac{(101-1)^{6n-3}-1}{101}\\~\\ \implies \frac{(-1)^{6n-3}-1}{101}\\~\\ \implies \frac{(-1)^{-3}-1}{101}\\~\\ \implies \frac{-2}{101}\\~\\ \implies \frac{99}{101}\\~\\ \implies 99 }\end{align}\)

ganeshie8 (ganeshie8):

Neat :) binomial theorem will do !

OpenStudy (anonymous):

Bravo :3

OpenStudy (mathmath333):

i was kind of stuck at -2 and forgot that it equals 99 ,lol

OpenStudy (anonymous):

oh :D the real challenge is to for the orginal problem that the mod is 1

OpenStudy (anonymous):

show*

hartnn (hartnn):

why do i feel, all of the @mathmath333 solution are similar :3 lol

OpenStudy (anonymous):

mathmath style :S

OpenStudy (anonymous):

well i have to say this for u hartnn when u dont understand things they will look like the same :P

OpenStudy (mathmath333):

cuz i didnt use the mod technique but i basically learnt that to find remainders quickly and i think it is optimal than it

ganeshie8 (ganeshie8):

he is biased to binomial thm like me and all other clever mathematicaians haha!

OpenStudy (mathmath333):

^

OpenStudy (anonymous):

awwww xD

OpenStudy (anonymous):

this is what i call applied :P

hartnn (hartnn):

binomial theory rocks mod technique s....nvm

OpenStudy (anonymous):

:3 what i have to say 3 indians vs me

OpenStudy (mathmath333):

" mod arithmatic is universal"-ganeshie

ganeshie8 (ganeshie8):

mod/congruence theory is introduced by Gauss and i feel it is just a neat way to use division algorithm i thyink thats the reason the idea was rejected initially by math community.. i heard of this interesting story somewhere on youtube..xD

OpenStudy (anonymous):

im Gaussian lover

ganeshie8 (ganeshie8):

yes division algorithm is general so mod arithmetic/congruences are also general they make it more generic in group theory i think..

OpenStudy (anonymous):

guys how about this idea? \[\Large 101 \in \phi(\frac{ 10^{12n-1}-1 }{ 99 })\] if we prove that then it gives us that 101 isn't a factor of that.

OpenStudy (anonymous):

but phi Euler is very difficult XD

ganeshie8 (ganeshie8):

lol Gaussian lover ! it would have been tough for me to choose betwen euler and gauss if they were girls.. :P Euler dude never rests. I think he shows up more frequently compared to gauss

OpenStudy (anonymous):

Just \(\Large Fermet\) and \( \Large Euler \) XD

ganeshie8 (ganeshie8):

@PFEH.1999 for \(\large 10^{12n-6}-1 \pmod {99\cdot 101}\), try finding the order of \(10\) in \(\mod 99 \cdot 101\) by first finding \(\phi(99\cdot 101)\) that works perfectly!

OpenStudy (anonymous):

nice ;) will work on that...

OpenStudy (anonymous):

well i love 10 per time :D so its not a problem :3

OpenStudy (anonymous):

oh lets see if phi works

ganeshie8 (ganeshie8):

it will work thats how i would work when my brain doesn't work >.<

ganeshie8 (ganeshie8):

i have used it in your previous problem to prove 17 does not divide this number, right ?

OpenStudy (anonymous):

10^phi(99.101)=1 mod 99.101 10^(6000)=1 mod (99.101) eh yeah too easy :|

ganeshie8 (ganeshie8):

:D

OpenStudy (anonymous):

i expected it to be so hard :D

OpenStudy (anonymous):

@ganeshie8 <3 :*

OpenStudy (anonymous):

13:13